[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Mon Sep 04, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Mon Sep 4 20:28:07 UTC 2017


HSU Shift Log for Mon Sep 04 2017 16:27:58 GMT-0400 (EDT)

Shift: H

Forecaster: D Brown

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues: G-IV mission taking place this afternoon. Special  
soundings beginning over the central US at 18z today. Spread eastward  
tomorrow. Mike has coordinated these with the SDM. Updated G-16 meso  
request for 21-06z today/tonight and also put one in for tonight/tomorrow  
06-18z.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170904/409bdbfe/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list